What is the volume of a cube that is 15' x 20' x 6"
PLUS
the volume of a cube that is 15' x 23' 6"
PLUS
HALF the volume of a cylinder that is 6" deep and has a radius of 7.5'

Answer by Alan3354(69443)

What is the volume of a cube that is 15' x 20' x 6"
15*20*0.5 = 150 cubic feet
= 150*1728 = 259200 cubic inches
----------------
PLUS
the volume of a cube that is 15' x 23' 6"
A cube has 3 dimensions, only 2 are given.
------------------
PLUS
HALF the volume of a cylinder that is 6" deep and has a radius of 7.5'
V = pi*r^2*h
V = pi*7.5^2*0.5/2 cubic feet
v =~44.1786 cubic ft
V =~ 76,340.7 cubic inches